# A perspective on inelastic light scattering spectroscopy for probing transport of collective acoustic excitations

### v1

@@ I.2  Spontaneous Brillouin scattering spectroscopy
In principle, Stokes and anti-Stokes scattering of spontaneous Brillouin scattering are driven by modulations in the dielectric constant due to thermal fluctuations and zero-point energy Damzen et al. (2003); Boyd et al. (2008). Such variations in the optical properties lead to inelastic scattering of the incident light, and two types of excitations are involved depending on the optical characteristics of the materials. In weakly absorbing materials such as optically transparent or semitransparent media, where their optical penetration depth is sufficiently larger than the wavelength of incident light, the BLS signal primarily arises from bulk phonons via opto-elastic coupling Bottani and Fioretto (2018). In opaque or strongly absorbing materials, if penetration depth is limited to the surface due to high optical extinction of the medium, scattering predominantly occurs by surface phonons Briggs (2013). In either case, the magnitude of the measured phonon wavevector depends on the incident angle and the wavelength of the excitation laser as described in Eq. 1.
@@
A schematic of a representative experimental setup used for backscattering BLS is shown in Fig. 2(a). A laser beam is initially directed toward a beam splitter, where the reflected light is used to define a zero-frequency position for calibrating 3FP interferometers. The transmitted beam is refined by a polarizer, then its polarization is adjusted by a half-wave plate to enhance sensitivity to the quasiparticle mode of interest Heiman et al. (1979). The sample is typically mounted on a motorized rotation stage to vary the incidence angle of the laser Link et al. (2006); Pang et al. (2022), which provides access to different phonon wavevectors in opaque samples, whereas in bulk transparent materials, phonon wavevector is determined by the refractive index of the medium as $\lvert\mathbf{q}\rvert=4\pi n/\lambda_{i}$ in backscattering geometry Mutti et al. (1995). The scattered light is then focused into FP interferometers, which provide a high spectral contrast to isolate weak Brillouin peaks from strong elastic Rayleigh line while suppressing parasitic transmission peaks and maintaining high effective transmittance via tandem design of interferometers Sandercock (1982). The output is detected by a photomultiplier tube in single-photon counting mode to capture low-intensity scattered light. With sub-GHz spectral resolution, modern BLS configuration can detect weak optical signals, which allows access to hypersonic phonons and thermal magnons Sandercock (1982); Birt et al. (2013), even in thin film with a few Angstroms Hillebrands et al. (1987). Another advantage of such configuration is automation via self-aligning FP interferometers Hillebrands (1999), allowing versatility in typical lab settings. Adding an objective lens with a high numerical aperture to achieve small laser spot and integrating an automated 3D scanning stage yield micro-focused BLS, or $\mu$-BLS, which has been used for spatiotemporal imaging of surface phonon and spin waves Geilen et al. (2020); Yang et al. (2025); Madami et al. (2012).
@@
Figure 2: Typical experimental setup for (a) spontaneous Brillouin light scattering spectroscopy, adapted with permission from DOI:10.1088/0022-3727/45/27/275302, J. Phys. D: Appl. Phys., 45, 275302. Copyright 2012 IOP Publishing Ltd Guerette and Huang (2012), and (b) impulsive stimulated scattering spectroscopy.
@@ II  Applications of BLS
Figure 3: (a) Comparison of Brillouin spectra for hybrid halide perovskites with different cations and halides compositions, showing larger Brillouin shifts with increasing MA and Cl content. Reproduced from Ref. Kabakova et al. (2018) with permission from the Royal Society of Chemistry. (b) Phonon dispersion relation in a holey Si membrane obtained from BLS compared with simulations, showing an omnidirectional band gap Florez et al. (2022). Reprinted with permission from Florez, O. et al. Nature Nanotechnol. 17, 947-951 (2022). Copyright 2022 Springer Nature. (c) Measured Brillouin spectra of GaAs nanowires array (red) compared to bulk GaAs substrate (blue), showing additional low-frequency peaks corresponding to confined acoustic phonon modes. Reproduced from Ref. Kargar et al. (2016), licensed under a Creative Commons Attribution 4.0 International License. (d) Wavenumber-resolved BLS signal from a YIG thin film, showing field positions of maximum BLS amplitude, which corresponds to hybridized magnon-phonon modes. Holanda et al. (2018). Adapted with permission from J. Holanda et al. Nature Phys. 14, 500-506 (2018). Copyright 2018 Springer Nature.
@@
Conventionally, one of the earliest and continuing applications of BLS is characterizing elastic properties such as sound velocities and elastic moduli of materials Grimsditch and Ramdas (1975); Vacher and Boyer (1972); Whitfield et al. (1976); Yamaguchi et al. (1999), particularly bulk crystals and minerals under high pressures up to $\sim 64$ GPa and temperatures above 1800 K Whitfield et al. (1976); Zha et al. (2000); Jiang et al. (2009); Raetz et al. (2019); Kurnosov et al. (2024), where traditional mechanical testing is challenging to apply and only feasible up to a few GPa due to large sample size constraints and difficulties in ensuring sufficient transducer-sample adhesion Brazhkin and Lyapin (2001); Grimsditch et al. (1986); Polian et al. (1982); Zouboulis et al. (1998).
@@
More recently, Brillouin spectroscopy has been employed to characterize elastic properties of novel materials, including soft lattices of hybrid halide perovskites Létoublon et al. (2016); Kabakova et al. (2018); as shown in variations in Brillouin shifts from hybrid perovskites with different ionic compositions in Fig. 3(a), it was demonstrated that adjusting cation and halide composition can stiffen their lattices Kabakova et al. (2018). BLS has also been applied to complex crystals Stevens and Eckhardt (2005); Tan et al. (2012); Sanchez-Valle et al. (2005); Dirras et al. (2016); Pang et al. (2022), where a recent study used Brillouin scattering to obtain an elastic tensor of a topological insulator (Sb2Te3), resolving its out-of-plane elastic constant $c_{33}$ by measuring velocity of surface acoustic modes Baranowski et al. (2025). Likewise, BLS has been applied to 2D layered crystalline materials and polycrystalline thin films to measure elastic anisotropy Faurie et al. (2010); Jimenez-Rioboo et al. (2018); Reed and Koski (2022); Reed et al. (2024). For instance, in few-layer MoSe2 membranes, BLS was used to detect a $\sim 30\%$ reduction in their in-plane stiffness compared to bulk, confirming elastic softening as the number of layers decreases Babacic et al. (2021). Extending its scope, theoretical work has proposed that Brillouin scattering could provide access to Majorana fermions and other low-energy boundary modes arising in quantum spin liquids, in particular $\alpha-\mathrm{RuCl}_{3}$ 2D honeycomb structure, where these excitations are inaccessible using standard experimental probes such as angle-resolved photoemission spectroscopy due to their chargeless nature Perreault et al. (2016).
@@
In addition to elastic property measurements, BLS has also become an indispensable method for probing dispersion relations and lifetimes of gigahertz vibrational modes in nanostructures and phononic crystals Balandin and Nika (2012); Zhang et al. (2012); Hou et al. (2014); Mielcarek et al. (2012); Graczykowski et al. (2012); Nataj et al. (2023); Huang et al. (2020); Florez et al. (2022); Wright et al. (2024). In periodic nanostructured solids, BLS can directly observe phenomena such as Bragg band gaps, zone folding of acoustic phonon branches, and localized resonant modes Hou et al. (2014); Mielcarek et al. (2012); Yudistira et al. (2016). It is generally accepted that in surface acoustic waves, periodically arranged scatterers (holey structure) cause phonon band folding and Bragg gaps Florez et al. (2022), as shown in Fig. 3(b), while periodic arrays of local resonator (pillar-based) structures introduce additional low-frequency gaps Graczykowski et al. (2015); Sledzinska et al. (2020). Another interesting phenomenon in nanostructures is phonon confinement induced by boundary conditions, leading to modified phonon dispersion that can affect thermal conductivity Balandin and Wang (1998); Kargar et al. (2015); Cuffe et al. (2012); Neogi et al. (2015); Reed et al. (2019). As demonstrated in Fig. 3(c), BLS was employed to observe multiple acoustic phonon subbands in free-standing GaAs nanowires, confirming the existence of confined phonon modes in structures with length scales on the order of hundred nanometer Kargar et al. (2016).
@@
Another class of quasiparticles accessible by BLS are spin waves, also known as magnons. In magnetic thin films and multilayer heterostructures, BLS is one of the established techniques to measure dispersion relations and damping of spin waves Sandercock (1975); Sandercock and Wettling (1978); Grimsditch et al. (1979); An et al. (2014); Stashkevich et al. (2015); Ma et al. (2018). BLS has been used to spatially map the intensity of the Brillouin peak due to magnons under external magnetic field, thereby measuring the decay length of spin waves in thin metal strips Vogt et al. (2009); Demidov et al. (2009); Demidov and Demokritov (2015). Magnetoelastic waves are another area of interest, in which BLS has been employed to probe magnon-phonon interactions Agrawal et al. (2013); Zhao et al. (2021); Holanda et al. (2021); Kunz et al. (2024), demonstrating magnon-phonon conversion that yields a plateau in the field position since phonon modes are independent of the external magnetic field, as shown in Fig. 3(d), and further revealing that spin angular momentum can be transferred from magnons to phonons Holanda et al. (2018). Analogous to phonons, magnon confinement effects were observed using BLS and $\mu$-BLS in metallic thin films Demokritov et al. (2001); Wang et al. (2023); Demidov et al. (2008); Kargar et al. (2020), highlighting possible strategies to manipulate magnon transport for spintronics applications.
@@ III  Working principles and instrumentation of ISS
Impulsive stimulated light scattering is a method that probes light scattering using impulsive pump-probe setups to detect the transport and relaxation processes of energy carriers such as phonons and electrons Eichler et al. (2013). The scattering can be stimulated when the acoustic wave in a material is induced through the excitation process with the incident light.
@@
The excitation process through the interaction of the laser and acoustic waves in a material is explained by two primary mechanisms: electrostriction and optical absorption. First, electrostriction is a mechanism in which mechanical stress in materials varies under an applied electric field, thereby enabling the stimulated scattering in non-absorptive materials Kroll (1965); Beugnot and Laude (2012). Second, photothermal excitation can occur in absorbing materials Herman and Gray (1967); Pohl and Kaiser (1970), where optical absorption leads to localized heating and subsequent thermal expansion Boyd et al. (2008). In general, the stimulated stress in a sample is generated by a combination of both electrostriction and absorptive effects, with the relative contribution of each mechanism depending on the properties of the material.
@@
In non-absorptive media, Brillouin scattering process can be stimulated through the process of electrostriction. Impulsive stimulated Brillouin scattering (ISBS) is an excitation mechanism that utilizes electrostriction to generate dynamic gratings. The interference pattern modulates dielectric properties at regions of high electric field intensity, inducing a periodic density grating through electrostriction Kroll (1965); Beugnot and Laude (2012). This density wave propagates at the material’s sound velocity, with stress profiles that can be described by an impulse response. The generated acoustic waves exhibit dependence on the material’s intrinsic viscoelastic properties through their velocity and attenuation characteristics Torchinsky et al. (2009).
@@
On the other hand, for absorptive media, impulsive stimulated thermal scattering (ISTS) can measure acoustic properties of materials from their response to impulsive photothermal excitation Rogers et al. (2012). Incident optical beams excite the electrons, which, subsequently thermalize through recombination and scattering processes with the lattice, imparting their energy to phonons. This non-radiative recombination process of excited electrons dissipates energy as heat through phonons and includes the relaxation process after the Auger and Shockley–Read–Hall (trap-assisted) recombination. As a result, following the intensity grating of incident beams, a spatially periodic transient temperature profile is formed Choudhry et al. (2021). The resulting thermal expansion produces a subsequent stress, which launches counter-propagating coherent acoustic waves Kim et al. (2017).
@@
As an experimental method for measuring scattered light, ISS can be utilized in the form of transient grating spectroscopy (TGS). TGS is a simplified variant of four-wave mixing that employs two temporally coincident, coherent pump beams crossed at an angle to generate a spatially modulated excitation field in a sample. The interference between these two coherently superimposed optical pump waves creates a spatially periodic interference pattern, which can then be measured by probe beams Choudhry et al. (2021). Through variations in the real and imaginary refractive indices $n$ and $k$, the interference pattern generates phase and amplitude transient gratings, respectively, and the detector observes the probe beam diffracted from these gratings.
@@
Contemporary TGS configuration that involves the phase mask and implements heterodyne detection is presented in Fig. 2(b). Heterodyne detection is a phase-sensitive detection technique that mixes the diffracted beam with a reference beam Maznev et al. (1998). It offers advantages in achieving a high signal-to-noise ratio by suppressing unwanted noise components, while not saturating the detector. The first-order diffracted beams from the phase mask are selected as probe beams. One beam then passes through an optical phase adjust and acts as the signal beam to detect material properties, while the other beam is attenuated and employed as the reference beam in heterodyne detection. After Bragg diffraction of the signal beam at the sample, the diffracted signal beam is now superimposed onto the reflected or transmitted reference beam, depending on whether reflection or transmission geometry is used. For a thin sample under the transmission geometry, the intensity of the first-order diffracted probe beam under direct (homodyne) detection, assuming small refractive index modulation, is given by Eichler et al. (2013)
@@
and $t_{r}$ is the signal-to-reference transmission coefficient. Equation 4 indicates that several advantages can be obtained through the implementation of heterodyne detection. First, ISS signal magnitude can be increased by the attenuation factor ($t_{r}$), thereby achieving a high signal-to-noise ratio. Second, contributions from the amplitude and phase gratings can be decoupled by adjusting the relative phase $\Delta\phi$ between the signal and reference beams to be multiples of $\pi/2$. More precisely, through this phase adjustment, heterodyne detection produces a signal that varies with the material response parameters $\Delta n$ or $\Delta k$ and enables separation of the amplitude grating contribution from $\Delta k$ and the phase grating contribution from $\Delta n$ Choudhry et al. (2021). Additionally, subtracting experimental results between different $\Delta\phi$ can filter out unwanted signal components while remaining signal is doubled. Using this transient thermal grating approach, TGS provides information about thermoelastic properties of materials, such as thermal conductivity and elastic constants, through thermoelastic response measurements across various research fields, with specific applications detailed in a later section.

### v2

@@ I.2  Spontaneous Brillouin scattering spectroscopy
In principle, Stokes and anti-Stokes scattering of spontaneous Brillouin scattering are driven by modulations in the dielectric constant due to thermal fluctuations and zero-point energy damzen_2003_stimulated; Boyd_2008_textbook. Such variations in the optical properties lead to inelastic scattering of the incident light, and two types of excitations are involved depending on the optical characteristics of the materials. In weakly absorbing materials such as optically transparent or semitransparent media, where their optical penetration depth is sufficiently larger than the wavelength of incident light, the BLS signal primarily arises from bulk phonons via opto-elastic coupling Bottani_APX_2018. In opaque or strongly absorbing materials, if penetration depth is limited to the surface due to high optical extinction of the medium, scattering predominantly occurs by surface phonons briggs_text_2013. In either case, the magnitude of the measured phonon wavevector depends on the incident angle and the wavelength of the excitation laser as described in Eq. 1.
@@
A schematic of a representative experimental setup used for backscattering BLS is shown in Fig. 2(a). A laser beam is initially directed toward a beam splitter, where the reflected light is used to define a zero-frequency position for calibrating 3FP interferometers Bencivenga_revsciinst_2012. The transmitted beam is refined by a polarizer, then its polarization is adjusted by a half-wave plate to enhance sensitivity to the quasiparticle mode of interest heiman_prb_1979. The sample is typically mounted on a motorized rotation stage to vary the incidence angle of the laser link_japhys_2006; pang_adphr_2022, which provides access to different phonon wavevectors in opaque samples, whereas in bulk transparent materials, phonon wavevector is determined by the refractive index of the medium as $\lvert\mathbf{q}\rvert=4\pi n/\lambda_{i}$ in backscattering geometry mutti_adv_1995. The scattered light is then focused into FP interferometers, which provide a high spectral contrast to isolate weak Brillouin peaks from strong elastic Rayleigh line while suppressing parasitic transmission peaks and maintaining high effective transmittance via tandem design of interferometers sandercock_book_1982; Blachowicz_1996_revsciinst; Dil_1981_applopt; Scarponi_prx_2017. The output is detected by a photomultiplier tube in single-photon counting mode to capture low-intensity scattered light Witschas_2010_applopt. With sub-GHz spectral resolution, modern BLS configuration can detect weak optical signals, which allows access to hypersonic phonons and thermal magnons sandercock_book_1982; Bencivenga_revsciinst_2012; Birt_2013_apl, even in thin film with a few Angstroms hillebrands_prb_1987. Another advantage of such configuration is automation via self-aligning FP interferometers hillebrands_revsciinst_1999, allowing versatility in typical lab settings. Adding an objective lens with a high numerical aperture to achieve small laser spot and integrating an automated 3D scanning stage yield micro-focused BLS, or $\mu$-BLS, which has been used for spatiotemporal imaging of surface phonon and spin waves geilen_apl_2020; yang_pra_2025; camley_solphys_2012.
@@
Figure 2: Typical experimental setup for (a) spontaneous Brillouin light scattering spectroscopy, adapted with permission from DOI:10.1088/0022-3727/45/27/275302, J. Phys. D: Appl. Phys., 45, 275302. Copyright 2012 IOP Publishing Ltd BLS_setup, and (b) impulsive stimulated scattering spectroscopy.
Figure 3: (a) Comparison of Brillouin spectra for hybrid halide perovskites with different cations and halides compositions, showing larger Brillouin shifts with increasing MA and Cl content. Reproduced from Ref. kabakova2018effect with permission from the Royal Society of Chemistry. (b) Phonon dispersion relation in a holey Si membrane obtained from BLS compared with simulations, showing an omnidirectional band gap Florez_2022_NatNanotech. Reprinted with permission from Florez, O. et al., Nature Nanotechnol. 17, 947-951 (2022). Copyright 2022 Springer Nature. (c) Measured Brillouin spectra of GaAs nanowires array (red) compared to bulk GaAs substrate (blue), showing additional low-frequency peaks corresponding to confined acoustic phonon modes. Reproduced from Ref. Kargar_2016_Natcom_BLSGaAs, licensed under a Creative Commons Attribution 4.0 International License. (d) Wavenumber-resolved BLS signal from a YIG thin film, showing field positions of maximum BLS amplitude, which corresponds to hybridized magnon-phonon modes Holanda_2018_natphys. Adapted with permission from J. Holanda et al., Nature Phys. 14, 500-506 (2018). Copyright 2018 Springer Nature.
@@ II  Applications of BLS
Conventionally, one of the earliest and continuing applications of BLS is characterizing elastic properties such as sound velocities and elastic moduli of materials grimsditch_prb_1975; vacher_prb_1972; whitfield1976elastic; yamaguchi_jap_1999, particularly bulk crystals and minerals under high pressures up to $\sim 64$ GPa and temperatures above 1800 K whitfield1976elastic; zha_pnas_2000; jiang_physearth_2009; raetz_prb_2019; Kurnosov_2024_physchem, where traditional mechanical testing is challenging to apply and only feasible up to a few GPa due to large sample size constraints and difficulties in ensuring sufficient transducer-sample adhesion Brazhkin_jexptheo_2001; grimsditch1986brillouin; polian_prb_1982; zouboulis_prb_1998.
@@
More recently, Brillouin spectroscopy has been employed to characterize elastic properties of novel materials, including soft lattices of hybrid halide perovskites letoublon2016elastic_BLSapplication1; kabakova2018effect; as shown in variations in Brillouin shifts from hybrid perovskites with different ionic compositions in Fig. 3(a), it was demonstrated that adjusting cation and halide composition can stiffen their lattices kabakova2018effect. BLS has also been applied to complex crystals stevens_jchemphys_2005; tan_prl_2012; sanchez-valle_jap_2005; dirras_matscieng_2016; pang_adphr_2022, where a recent study used Brillouin scattering to obtain an elastic tensor of a topological insulator (Sb2Te3), resolving its out-of-plane elastic constant $c_{33}$ by measuring velocity of surface acoustic modes Baranowski_2025_scirep. Likewise, BLS has been applied to 2D layered crystalline materials and polycrystalline thin films to measure elastic anisotropy faurie_actmat_2010; jimenez-rioboo_apl_2018; reed_jap_2022; reed_acsnano_2024. For instance, in few-layer MoSe2 membranes, BLS was used to detect a $\sim 30\%$ reduction in their in-plane stiffness compared to bulk, confirming elastic softening as the number of layers decreases Babacic_2021_advmat. Extending its scope, theoretical work has proposed that Brillouin scattering could provide access to Majorana fermions and other low-energy boundary modes arising in quantum spin liquids, in particular $\alpha-\mathrm{RuCl}_{3}$ 2D honeycomb structure, where these excitations are inaccessible using standard experimental probes such as angle-resolved photoemission spectroscopy due to their chargeless nature Perreault_PRB_2016_MFermions.
@@
In addition to elastic property measurements, BLS has also become an indispensable method for probing dispersion relations and lifetimes of gigahertz vibrational modes in nanostructures and phononic crystals balandin2012phononics; Zhang_2012_apl; Hou_2014_apl; Mielcarek_2012_pss; Graczykowski_2012_prb; nataj_2023_natcomm; Huang_2020_nanotech; Florez_2022_NatNanotech; wright_apl_2024. In periodic nanostructured solids, BLS can directly observe phenomena such as Bragg band gaps, zone folding of acoustic phonon branches, and localized resonant modes Hou_2014_apl; Mielcarek_2012_pss; Yudistira_2016_prb. It is generally accepted that in surface acoustic waves, periodically arranged scatterers (holey structure) cause phonon band folding and Bragg gaps Florez_2022_NatNanotech, as shown in Fig. 3(b), while periodic arrays of local resonator (pillar-based) structures introduce additional low-frequency gaps Graczykowski_PRB_2015; Sledzinska_2020_advfuncmat. Another interesting phenomenon in nanostructures is phonon confinement induced by boundary conditions, leading to modified phonon dispersion that can affect thermal conductivity balandin_prb_1998; kargar_apl_2015; cuffe_nanolett_2012; neogi_acsnano_2015; reed_nanolett_2019. As demonstrated in Fig. 3(c), BLS was employed to observe multiple acoustic phonon subbands in free-standing GaAs nanowires, confirming the existence of confined phonon modes in structures with length scales on the order of hundred nanometer Kargar_2016_Natcom_BLSGaAs.
@@
Another class of quasiparticles accessible by BLS are spin waves, also known as magnons. In magnetic thin films and multilayer heterostructures, BLS is one of the established techniques to measure dispersion relations and damping of spin waves sandercock2007some; Sandercock_1978_ieee; Grimsditch_1979_prl; an_prb_2014; stashkevich_prb_2015; ma_prl_2018. BLS has been used to spatially map the intensity of the Brillouin peak due to magnons under external magnetic field, thereby measuring the decay length of spin waves in thin metal strips vogt_apl_2009; Demidov_2009_prl; demidov_ieee_2015. Magnetoelastic waves are another area of interest, in which BLS has been employed to probe magnon-phonon interactions agrawal_prl_2013; zhao_physrevappl_2021; holanda_apl_2021; kunz_apl_2024, demonstrating magnon-phonon conversion that yields a plateau in the field position since phonon modes are independent of the external magnetic field, as shown in Fig. 3(d), and further revealing that spin angular momentum can be transferred from magnons to phonons Holanda_2018_natphys. Analogous to phonons, magnon confinement effects were observed using BLS and $\mu$-BLS in metallic thin films Demokritove_2001_physrep; wang_prx_2023; demidov_apl_2008; Kargar_2020_jmagmag, highlighting possible strategies to manipulate magnon transport for spintronics applications.
@@ III  Working principles and instrumentation of ISS
Impulsive stimulated light scattering is a method that probes light scattering using impulsive pump-probe setups to detect the transport and relaxation processes of energy carriers such as phonons and electrons Eichler_2013_textbook. The scattering can be stimulated when the acoustic wave in a material is induced through the excitation process with the incident light.
@@
The excitation process through the interaction of the laser and acoustic waves in a material is explained by two primary mechanisms: electrostriction and optical absorption. First, electrostriction is a mechanism in which mechanical stress in materials varies under an applied electric field, thereby enabling the stimulated scattering in non-absorptive materials Kroll_JAP_1965; Beugnot_PRB_2012. Second, photothermal excitation can occur in absorbing materials Herman_PRL_1967; Kaiser_PRB_1970, where optical absorption leads to localized heating and subsequent thermal expansion Boyd_2008_textbook. In general, the stimulated stress in a sample is generated by a combination of both electrostriction and absorptive effects, with the relative contribution of each mechanism depending on the properties of the material.
@@
In non-absorptive media, Brillouin scattering process can be stimulated through the process of electrostriction. ISBS is an excitation mechanism that utilizes electrostriction to generate dynamic gratings. The interference pattern modulates dielectric properties at regions of high electric field intensity, inducing a periodic density grating through electrostriction Kroll_JAP_1965; Beugnot_PRB_2012. This density wave propagates at the material’s sound velocity, with stress profiles that can be described by an impulse response. The generated acoustic waves exhibit dependence on the material’s intrinsic viscoelastic properties through their velocity and attenuation characteristics Torchinsky_JCP_2009.
@@
On the other hand, for absorptive media, ISTS can measure acoustic properties of materials from their response to impulsive photothermal excitation Rogers_2002_textbook_ISTS. Incident optical beams excite the electrons, which, subsequently thermalize through recombination and scattering processes with the lattice, imparting their energy to phonons. This non-radiative recombination process of excited electrons dissipates energy as heat through phonons and includes the relaxation process after the Auger and Shockley–Read–Hall (trap-assisted) recombination. As a result, following the intensity grating of incident beams, a spatially periodic transient temperature profile is formed Choudhry_2021_JAP. The resulting thermal expansion produces a subsequent stress, which launches counter-propagating coherent acoustic waves kim_aplmat_2017.
@@
As an experimental method for measuring scattered light, ISS can be utilized in the form of transient grating spectroscopy (TGS). TGS is a simplified variant of four-wave mixing that employs two temporally coincident, coherent pump beams crossed at an angle to generate a spatially modulated excitation field in a sample. The interference between these two coherently superimposed optical pump waves creates a spatially periodic interference pattern, which can then be measured by probe beams Johnson_2012_JAP; vegaflick_revsciins_2015; Dennett_JAP_2018; Choudhry_2021_JAP. Through variations in the real and imaginary refractive indices $n$ and $k$, the interference pattern generates phase and amplitude transient gratings, respectively, and the detector observes the probe beam diffracted from these gratings.
@@
Contemporary TGS configuration that involves the phase mask and implements heterodyne detection is presented in Fig. 2(b). Heterodyne detection is a phase-sensitive detection technique that mixes the diffracted beam with a reference beam Maznev_1998_OptLett; Goodno_Opt_1998; Torre_PRE_2001; Taschin_PRE_2006. It offers advantages in achieving a high signal-to-noise ratio by suppressing unwanted noise components, while not saturating the detector. The first-order diffracted beams from the phase mask are selected as probe beams. One beam then passes through an optical phase adjust and acts as the signal beam to detect material properties, while the other beam is attenuated and employed as the reference beam in heterodyne detection. After Bragg diffraction of the signal beam at the sample, the diffracted signal beam is now superimposed onto the reflected or transmitted reference beam, depending on whether reflection or transmission geometry is used. For a thin sample under the transmission geometry, the intensity of the first-order diffracted probe beam under direct (homodyne) detection, assuming small refractive index modulation, is given by Eichler_2013_textbook
@@
and $t_{r}$ is the signal-to-reference transmission coefficient. Equation 4 indicates that several advantages can be obtained through the implementation of heterodyne detection. First, scattered signal magnitude can be increased by the attenuation factor ($t_{r}$), thereby achieving a high signal-to-noise ratio. Second, contributions from the amplitude and phase gratings can be decoupled by adjusting the relative phase $\Delta\phi$ between the signal and reference beams to be multiples of $\pi/2$. More precisely, through this phase adjustment, heterodyne detection produces a signal that varies with the material response parameters $\Delta n$ or $\Delta k$ and enables separation of the amplitude grating contribution from $\Delta k$ and the phase grating contribution from $\Delta n$ Choudhry_2021_JAP. Additionally, subtracting experimental results between different $\Delta\phi$ can filter out unwanted signal components while remaining signal is doubled. Using this transient thermal grating approach, TGS provides information about thermoelastic properties of materials, such as thermal conductivity and elastic constants, through thermoelastic response measurements across various research fields, with specific applications detailed in a later section.
Figure 4: (a) Measured (solid circles) and calculated (line) dispersion relation of the lowest frequency surface acoustic mode in a silicon thin film (Inset: dispersion curves of the four lowest frequency modes) maznev_apl_2009. Reprinted from A. A. Maznev et al., Appl. Phys. Lett. 95, 011903 (2009), with the permission of AIP Publishing. (b) TGS signal from a micro-patterned 1D phononic structure, showing acoustic oscillations associated with non-leaky, long-lived surface acoustic modes. (Inset: the same signal over a longer time scale) maznev_japplphys_2009. Reprinted from A. A. Maznev et al., J. Appl. Phys. 105, 123530 (2009), with the permission of AIP Publishing. (c) Acoustic dispersion of a 2D granular crystal, where red and black markers represent the measured frequencies with and without silica microspheres, respectively. A horizontal dotted line indicates the resonance frequency arising from microsphere-substrate adhesion boechler2013interaction. Reprinted figure with permission from N. Boechler et al., Interaction of a Contact Resonance of Microspheres with Surface Acoustic Waves, Phys. Rev. Lett. 111, 036103 (2013). Copyright 2013 by the American Physical Society. (d) Measured spin helix lifetimes versus wavevector in GaAs/AlGaAs quantum wells, showing enhanced lifetime peaks to higher wavevectors with increasing doping asymmetry koralek2009emergence. Reproduced with permission from J. D. Koralek et al., Emergence of the persistent spin helix in semiconductor quantum wells, Nature 458, 7238 (2009). Copyright 2009 Springer Nature.